# Active Gel Theory for Cell Migration With Two Myosin Isoforms

### v1

@@ Abstract
Myosin II molecular motors slide actin filaments relatively to each other and are essential for force generation, motility and mechanosensing in animal cells. For non-muscle cells, evolution has resulted in three different isoforms, which have different properties concerning the motor cycle and also occur in different abundances in the cells, but their respective biological and physical roles are not fully understood. Here we use active gel theory to demonstrate the complementary roles of isoforms A and B for cell migration. We first show that our model can be derived both from coarse-graining kinetic equations and from nonequilibrium thermodynamics as the macroscopic limit of a two-component Tonks gas. We then parametrize the model and show that motile solutions exist, in which the more abundant and more dynamic isoform A is localized to the front and the stronger isoform B to the rear, in agreement with experiments. We then explore the parameter space of the model and find a general pull-and-push mechanism that can produce different migratory modes, including cell oscillations in length and velocity. We also describe an analytical solution for the stiff limit. Our findings highlight the importance of including isoform-specific molecular details to describe whole cell behavior.
@@ I Introduction
Cell migration is an essential process for any living organism. In multicellular animals like humans, it is important mainly for embryonic development [1], wound healing [2] and the immune response [3], but also for the spread of cancer [4]. It is strongly linked to the underlying molecular processes, because cells have to generate force and movement that is converted into consistent migration on the cell scale. This function is provided by the cytoskeleton, a collection of polymer networks. Cells can push their envelope forward by polymerization of these filaments. They also can generate pulling forces by sliding them relative to each other with molecular motors. For animal cells, pushing and pulling forces are generated mainly by actin filaments and myosin II molecular motors. In cell migration, they go hand-in-hand, because the actin networks generated by polymerization are pulled backwards by myosin II motors. The resulting retrograde flow is then transmitted to the substrate by adhesions and thus generates forward motion of the cell [5]. Migration requires a cyclic process that converts energy provided in the form of ATP into force and movement. Moreover, initiation of motility requires a symmetry break between front and back, which can be either spontaneous or driven by some external cue [6, 7, 8]. Usually this break of symmetry is related to activation of the Rac/Cdc42 and RhoA pathways, which for quantitative experiments can be controlled by optogenetics [9, 10, 11].
@@
In non-muscle animal cells contractile stress for cell motility is produced by motor proteins of the non-muscle myosin II (NMII) family [13]. Evolution has resulted in three isoforms, namely NMIIA, NMIIB and NMIIC [14]. After activation of the Rho-pathway, the myosin motors are assembled into mini-filaments that typically mix different isoforms [15, 14]. These mini-filaments then bind to the actin cytoskeleton, generate stress and at the same time are advected with the resulting cytoskeletal flow. In the case of asymmetric contractile stress distributions, the cell is polarized and starts migrating. While NMIIA is most abundant and propels loads the quickest along actin filaments, NMIIB is the strongest due to its higher duty ratio; NMIIC seems to have little relevance for cell migration, also due to its low abundance [16, 12]. Recent experiments have shown that in a migrating cell the isoforms A and B are not distributed equally [15, 12]. While the more abundant and more dynamic isoform A localizes more at the leading edge, the stronger isoform B accumulates towards the trailing edge (cf. Fig. 1).
@@
The paper is organized as follows. We will start by defining the mathematical model based on active gel theory introducing the constitutive relation for stress. From binding kinetics we derive an effective nonlinear diffusion and demonstrate its consistency with the thermodynamics of a Tonks gas of two species. We then investigate the resulting closed boundary value problem and recapture experimental results in steady-state. Lifting the constraint of steady-state we demonstrate an oscillation mechanism for cell length and velocity. Lastly, we introduce a limiting case which can be treated analytically and for which we present a phase diagram indicating four different modes of motility.
@@ II.1 Geometry and Mechanics of the cell
Our one-dimensional model considers an infinitely compressible active gel. As commonly assumed [20, 22], the active contractile stress depends linearly on the motor concentration $c$, $\sigma_{act}=-\chi c$, resulting in the constitutive relation
@@
where $\xi$ is a friction coefficient. The cell occupies the range $x\in[l_{-},l_{+}]$ and hence its length is given by $L(t)=l_{+}(t)-l_{-}(t)$ and its midpoint velocity by $\dot{G}(t)=(v(l_{+},t)+v(l_{-},t))/2$. We use an elastic boundary condition, $\sigma(l_{\pm},t)=-k(L(t)-L_{0})/L_{0}$, that models membrane tension and all other restoring forces limiting the cell size [20, 21, 22]. Here $k$ is the effective spring constant of the cell and $L_{0}$ its homeostatic length.
@@ II.3 Myosin as Tonks gas
Within nonequilibrium thermodynamics, fluxes are driven by gradients in conjugate thermodynamic forces. The relevant couple in our case is particle flux and chemical potential. In linear theory, the particle flux is expanded to linear order and can be associated with the diffusion within the multicomponent mixture. For the gas of two species we obtain (in one spatial dimension) the relation
@@ III.1 Comparison to One-Species Models
As a background for the new two-species model, we briefly recall the basic features of the one-species model. As studied in [19, 20], the model for one species and a constant diffusion coefficient allows for a supercritical pitchfork bifurcation from sessile to moving cells as a function of Pe. The instability is due to the motors being advected to and accumulating at the back, breaking the symmetry. The fact that the motor distribution is strongly peaked at the trailing end motivated the addition of excluded volume in Ref. [22], implying nonlinear diffusion (namely an increase of diffusion close to crowding) which smoothens the motor distribution. There, it was also shown that attractive interactions, modelling myosin mini-filament assembly, can render the bifurcation from supercritical to subcritical, thus leading to bistability and the possibility of optogenetic switching.
@@ III.2 Model Parameters
Informed by experimental data on crawling cells, we first estimate the parameters of our model. Due to $\xi\sim 2\times 10^{14}Pa\ s\ m^{-1}$ [22, 41], $\eta\sim 10^{5}Pa\ s$ [20, 41], $k\sim 10^{4}Pa$ [42, 43, 16] and $L_{0}\sim 20\mu m$ [20] we obtain a (squared) relative viscous timescale of $\mathcal{L}^{2}=1.25$. Using $D_{\text{A}}\sim 0.1\times 10^{-12}m^{2}\ s^{-1}$ [44, 45, 46, 47], we infer the order of magnitude of the Péclet number to be $\mbox{{Pe}}\sim 100$.
The isoform-specific (off-)binding rates determine $K_{r}$, which in turn serves as a measure of relative diffusion. While the binding rates of both isoforms are equal ($k_{\text{A,on}}=k_{\text{B,on}}=0.2s^{-1}$), off-binding rates are $k_{\text{A,off}}=1.71s^{-1}$ and $k_{\text{B,off}}=0.35s^{-1}$ [28, 27, 48, 49] leading to a relative diffusion of $K_{r}\approx 0.2$. The contractility parameters, as defined earlier, are of the order $P_{i}\sim 0.1$ [20]. Due to the different duty ratios of the isoforms A and B we find $P_{\text{B}}\approx 3.6\cdot P_{\text{A}}$ [27] and, thus, choose $P_{\text{A}}=0.05$ and $P_{\text{B}}=0.18$. Lastly, following arguments from Ref. [22] we use $c_{\infty}=4$ for the saturation concentration. The concentration ratio of $c_{r}\approx 1.5$ is taken from experiments [12].
@@ III.3 Myosin Isoform Distributions in Steadily Moving Cells
At the critical Péclet number, a supercritical pitchfork bifurcation occurs, where the sessile state becomes unstable and stable motile steady states emerge. The two motile branches in panel (a) are $\pm$-symmetric and belong to the same, slightly contracted branch of the cell’s length in panel (b). Whether one moves to the right or to the left is determined by the initial conditions. In the steadily moving cell state beyond the bifurcation, the myosin distributions are asymmetric, as depicted in panel (e). This symmetry break in the concentrations induces a polarized active stress profile that initiates cell polarization and motility. In the shown example, the active stress is highest at the right edge ($u=1$), causing a negative velocity of the cell. The trailing edge is the locus of highest myosin concentration and, thus, highest active stress, in agreement to earlier theoretical works [20, 21] as well as experimental findings [12, 14, 16].
Figure 3: Oscillatory dynamics of the two isoform system. (a) Hopf bifurcation at $\mathcal{K}_{crit}=11500$ in the $\dot{G}_{max}$-Pe-plane, where $\dot{G}_{max}$ refers to the maximum velocity within an oscillation cycle. (b) Exemplary limit cycles in the $\dot{G}_{max}$-$\Delta c_{B}$-plane, where $\Delta c_{B}=c_{B}(l_{+})-c_{B}(l_{-})$ is a measure of the polarity of the myosin B distribution. The size of the limit cycle grows with Pe. (c) Exemplary limit cycles in the plane cell velocity vs. cell length. (d) Examples of the three different motility modes – sessile, transient oscillation and stable oscillation. Shown is the cell velocity over time. (e) Kymographs visualizing the pull-push-mechanism. Shown are the distributions of both myosin isoforms (upper and middle panel) and the resulting total active stress (lower panel) in a stable oscillatory state at $Pe=16000$. Parameters as in Fig. 2, except for concentration ratio of $c_{r}=1$ and strongly increased dissociation/diffusion ratio $K_{r}=10$.
@@
As already discussed, the two myosin isoforms and hence the concentration ratio are homogeneously distributed within the sessile state, see Fig. 2 (d). Comparing to Fig. 1 (d), due to the lack of a polarization axis in the sessile state, we should identify the lamellum with the cell center and the region called ”intermediate” region in Fig. 1 (d), and both edges with the periphery. In the motile steady state the stronger isoform B sits in the back, while the more diffusive isoform A is displaced towards the front causing the varying concentration ratios as shown in panel (f). This redistribution is due to the retrograde flow, which is towards the trailing edge and during which motors unbind from actin. The higher the diffusion, the less motors of the respective isoform arrive at the trailing edge. For the estimated value of $K_{r}=0.2$, isoform A has a larger effective diffusion than isoform B. Thus, B accumulates more quickly at the trailing edge. The effect of excluded volume then constrains A to accumulate further up front. The obtained results capture the experimental data from Weissenbruch et al. [12] qualitatively very well, compare Fig. 2(f) and Fig. 1(c).
@@ III.4 Exploring Parameter Space and Occurrence of Oscillations
We now lift the constraint of steady states, solving the full boundary value problem Eq. (10) numerically, and explore the parameter space beyond the values estimated from experiments. The most interesting parameter to study is the ratio of the dissociation constants, $K_{r}$, which directly translates into the ratio of the diffusion constants of the two isoforms, cf. Eq. (5). The case discussed in the previous section corresponded to a low $K_{r}=0.2$. We now consider a large $K_{r}=10$ and fix the concentration ratio to $c_{r}=1$ to focus on the effect of the reaction constants.
In this parameter regime, we find additional modes of motility, in particular the possibility of oscillations, as shown in Fig. 3. As before, for small Péclet number Pe we find stable sessile steady states. Increasing Pe leads to the onset of transient oscillations, whose amplitudes decay back to the sessile steady state in the regime $6000\lesssim Pe\lesssim 11500$. Beyond a critical value of $Pe_{\text{crit}}\approx 11500$, these oscillations prevail and become stable. This behavior can be traced back to a Hopf bifurcation as shown in Fig. 3(a)-(c). Examples for the three different modes of motility – sessile, transient oscillation and stable oscillation – are shown in Fig. 3 (d).
@@
How can the mechanism of the oscillation be understood? One has to note that for $K_{r}$ large, one myosin isoform – namely myosin B – is both more contractile and more diffusive. If we assume initially a spatial separation of both isoforms, the stronger isoform B then induces a larger – and hence dominating – contractile stress that induces an actin flow towards the position where isoform B is the most aggregated. Myosin motors of isoform A are advected along the retrograde actin flow. As an effect of volume exclusion, isoform A subsequently displaces the more diffusive isoform B from its initial position and effectively ”pushes” it to the other side of the cell. From this point on the pull-and-push cycle repeats. A schematic sketch of the internal myosin dynamics during one oscillation cycle is depicted in Fig. 4.

### v2

@@ Abstract
Myosin II molecular motors slide actin filaments relatively to each other and are essential for force generation, motility and mechanosensing in animal cells. For non-muscle cells, evolution has resulted in three different isoforms, which have different properties concerning the motor cycle and also occur in different abundances in the cells, but their respective biological and physical roles are not fully understood. Here we use active gel theory to demonstrate the complementary roles of isoforms A and B for cell migration. We first show that our model can be derived both from coarse-graining kinetic equations and from nonequilibrium thermodynamics as the macroscopic limit of a two-component Tonks gas. We then parametrize the model and show that motile solutions exist, in which the more abundant and more dynamic isoform A is localized to the front and the stronger isoform B to the rear, in agreement with experiments. Exploring parameter space beyond the isoform parameters typical for animal cells, we also find cell oscillations in length and velocity, which might be realized for genetically engineered systems. We also describe an analytical solution for the stiff limit, which then is used to calculate a state diagram, and the effect of actin polymerization at the boundaries, that leads to an imperfect pitchfork bifurcation. Our findings highlight the importance of including isoform-specific molecular details to describe whole cell behavior.
@@ I Introduction
Cell migration is an essential process for any living organism. In multicellular animals like humans, it is important mainly for embryonic development [1], wound healing [2] and the immune response [3], but also for the spread of cancer [4]. It is strongly linked to the underlying molecular processes, because cells have to generate force and movement that is converted into consistent migration on the cell scale. This function is provided by the cytoskeleton, a collection of polymer networks. Cells can push their envelope forward by polymerization of these filaments. They also can generate pulling forces by sliding them relative to each other with molecular motors. For animal cells, pushing and pulling forces are generated mainly by actin filaments and myosin II molecular motors. In cell migration, they go hand-in-hand, because the actin networks generated by polymerization are pulled backward by myosin II motors. The resulting retrograde flow is then transmitted to the substrate by adhesions and thus generates forward motion of the cell [5]. Migration requires a cyclic process that converts energy provided in the form of ATP into force and movement. Moreover, initiation of motility requires a symmetry break between front and back, which can be either spontaneous or driven by some external cue [6, 7, 8]. Usually this break of symmetry is related to activation of the Rac/Cdc42 and RhoA pathways, which for quantitative experiments can be controlled by optogenetics [9, 10, 11].
@@
In non-muscle animal cells contractile stress for cell motility is produced by motor proteins of the non-muscle myosin II (NMII) family [13]. Evolution has resulted in three isoforms, namely NMIIA, NMIIB and NMIIC [14]. After activation of the Rho-pathway, the myosin motors are assembled into minifilaments that typically mix different isoforms [15, 14]. These minifilaments then bind to the actin cytoskeleton, generate stress and at the same time are advected with the resulting cytoskeletal flow. In the case of asymmetric contractile stress distributions, the cell is polarized and starts migrating. While NMIIA is most abundant and propels loads the quickest along actin filaments, NMIIB is the strongest due to its higher duty ratio; NMIIC seems to have little relevance for cell migration, also due to its low abundance [16, 12]. Recent experiments have shown that in a migrating cell the isoforms A and B are not distributed equally [15, 12]. While the more abundant and more dynamic isoform A localizes more at the leading edge, the stronger isoform B accumulates towards the trailing edge (cf. Fig. 1).
@@
The paper is organized as follows. We will start by defining the mathematical model based on active gel theory introducing the constitutive relation for stress. From binding kinetics we derive an effective nonlinear diffusion and demonstrate its consistency with the thermodynamics of a Tonks gas of two species. We then investigate the resulting closed boundary value problem and recapture experimental results in steady-state. Lifting the constraint of steady-state we demonstrate an oscillation mechanism for cell length and velocity. Next we introduce a limiting case which can be treated analytically and for which we present a phase diagram indicating four different modes of motility. Lastly, we demonstrate the effect of actin polymerization.
@@ II.1 Geometry and Mechanics of the cell
Our one-dimensional model considers an infinitely compressible active gel. As commonly assumed [24, 20, 22], the active contractile stress depends linearly on the motor concentration $c$, $\sigma_{act}=-\chi c$, resulting in the constitutive relation
@@
where $\xi$ is a friction coefficient. The cell occupies the range $x\in[l_{-},l_{+}]$ and hence its length is given by $L(t)=l_{+}(t)-l_{-}(t)$ and its midpoint velocity by $\dot{G}(t)=\frac{d}{dt}(l_{+}(t)+l_{-}(t))/2$. We use an elastic boundary condition, $\sigma(l_{\pm},t)=-k(L(t)-L_{0})/L_{0}$, that models membrane tension and all other restoring forces limiting the cell size [20, 21, 22]. Here $k$ is the effective spring constant of the cell and $L_{0}$ its homeostatic length.
@@ II.3 Myosin as Tonks gas
Within nonequilibrium thermodynamics, fluxes are driven by gradients in conjugate thermodynamic forces. The relevant couple in our case is particle flux and chemical potential. We use standard linear theory to express particle fluxes as a linear expansion of all thermodynamic forces, introducing the phenomenological Onsager coefficients $L_{ij}$ that in turn can also depend on the thermodynamic variables. Doing so we can relate the particle flux to the diffusion within the multicomponent mixture
@@ III.1 Comparison to One-Species Models
As a background for the new two-species model, we briefly recall the basic features of the one-species model. As studied in [19, 20], the model for one species and a constant diffusion coefficient allows for a supercritical pitchfork bifurcation from sessile to moving cells as a function of Pe. The instability is due to the motors being advected to the back by retrograde flow, from where they pull further motors backward in a positive feedback loop. Because the motor distribution at the back can become unrealistically high in this model, excluded volume effects have been introduced, which imply nonlinear diffusion and smoothen the motor distribution [22]. It was also shown that attractive interactions, modeling myosin minifilament assembly, can render the bifurcation subcritical, thus leading to bistability and the possibility of optogenetic switching. More recently, it has been shown in a mathematically rigorous analysis that this subcritical bifurcation persists in a 2D version of this model [43].
@@ III.2 Model Parameters
Informed by experimental data on crawling cells, we first estimate the parameters of our model. Due to $\xi\sim 2\times 10^{14}\,\mathrm{Pa}\,\mathrm{s}\,\mathrm{m}^{-1}$ [22, 44], $\eta\sim 10^{5}\mathrm{Pa}\,\mathrm{s}$ [20, 44], $k\sim 10^{4}\,\mathrm{Pa}$ [45, 46, 16] and $L_{0}\sim 20\,\mu\mathrm{m}$ [20] we obtain a (squared) relative viscous length scale of $\mathcal{L}^{2}=1.25$. Using $D_{\text{A}}\sim 0.1\times 10^{-12}\,\mathrm{m}^{2}\,\mathrm{s}^{-1}$ [47, 48, 49, 50], we infer the order of magnitude of the Péclet number to be $\mbox{{Pe}}\sim 100$.
The isoform-specific (off-)binding rates determine $K_{r}$, which in turn serves as a measure of relative diffusion. While the binding rates of both isoforms are equal ($k_{\text{A,on}}=k_{\text{B,on}}=0.2\,\mathrm{s}^{-1}$), off-binding rates are $k_{\text{A,off}}=1.71\,\mathrm{s}^{-1}$ and $k_{\text{B,off}}=0.35\,\mathrm{s}^{-1}$ [29, 28, 51, 52] leading to a relative diffusion of $K_{r}\approx 0.2$. The contractility parameters, as defined earlier, are of the order $P_{i}\sim 0.1$ [20]. Due to the different duty ratios of the isoforms A and B we find $P_{\text{B}}\approx 3.6\cdot P_{\text{A}}$ [28] and, thus, choose $P_{\text{A}}=0.05$ and $P_{\text{B}}=0.18$. Lastly, following arguments from Ref. [22] we use $c_{\infty}=4$ for the saturation concentration. The concentration ratio of $c_{r}\approx 1.5$ is taken from experiments [12].
@@ III.3 Myosin Isoform Distributions in Steadily Moving Cells
At the critical Péclet number, a supercritical pitchfork bifurcation occurs, where the sessile state becomes unstable and stable motile steady states emerge. The two motile branches in panel (a) are $\pm$-symmetric and belong to the same, slightly contracted branch of the cell’s length in panel (b). Whether one moves to the right or to the left is determined by the initial conditions. Due to the high dimensionality of our nonlinear PDE system, analytical proof of attractor uniqueness is infeasible. However, Recho et al. [20] showed that in a similar single-species model, only the first supercritical pitchfork bifurcation leads to stable modes. Given this similarity, we expect the same here. Numerical simulations from varied initial conditions revealed no evidence of additional stable branches.
@@
In the steadily moving cell state beyond the bifurcation, the myosin distributions are asymmetric, as depicted in panel (e). This symmetry break in the concentrations induces a polarized active stress profile that initiates cell polarization and motility. In the shown example, the active stress is highest at the right edge ($u=1$), inducing a retrograde flow of actin to this end. Therefore, the cell moves to the left, like in the experimental examples from Fig. 1. The trailing edge is the locus of highest myosin concentration and, thus, highest active stress, in agreement to earlier theoretical works [20, 21] as well as experimental findings [12, 14, 16].
@@
As already discussed, the two myosin isoforms and hence the concentration ratio are homogeneously distributed within the sessile state, see Fig. 2 (d). Comparing to Fig. 1 (d), due to the lack of a polarization axis in the sessile state, we should identify the lamellum with the cell center and the region called ”intermediate” region in Fig. 1 (d), and both edges with the periphery. In the motile steady state the stronger isoform B sits in the back, while the more diffusive isoform A is displaced towards the front causing the varying concentration ratios as shown in panel (f). This redistribution is due to the balance of retrograde flow and diffusion, the former from flow towards the trailing edge and the latter effected by motors unbinding from actin. The higher the diffusion, the less motors of the respective isoform arrive at the trailing edge. For the estimated value of $K_{r}=0.2$, isoform A has a larger effective diffusion than isoform B. Thus, B accumulates more quickly at the trailing edge. The effect of excluded volume then constrains A to accumulate further up front.
Figure 3: Kymograph, i.e. trajectories in space and time, showing the advection through retrograde actin flow of cell that approaches steady-state motility. The flow lines (red) are directed towards the rear end and the flow velocity reaches a maximum value of $\tilde{v}_{\text{max}}\sim 4.0\times 10^{-2}\mu\mathrm{m}\,\mathrm{s}^{-1}$.
The symmetry breaking of concentration profiles is caused by the advection of myosin along the actin retrograde flow. The diffusion, depending on both $K_{r}$ and $c_{\infty}$, counteracts this by flattening the profiles. Hence, this polarization can only occur if the Péclet number is sufficiently large. This mechanism remains qualitatively unchanged when varying the other parameters. A more thorough sensitivity analysis of the parameters is performed in Appendix A. The theoretical results agree with the experimental data from Weissenbruch et al. [12], compare Fig. 2(f) and Fig. 1(c).
Figure 4: Oscillatory dynamics of the two isoform system. (a) Hopf bifurcation at $\mathcal{K}_{crit}=11500$ in the $\dot{G}_{max}$-Pe-plane, where $\dot{G}_{max}$ refers to the maximum velocity within an oscillation cycle. (b) Exemplary limit cycles in the $\dot{G}_{max}$-$\Delta c_{B}$-plane, where $\Delta c_{B}=c_{B}(l_{+})-c_{B}(l_{-})$ is a measure of the polarity of the myosin B distribution. The size of the limit cycle grows with Pe. (c) Exemplary limit cycles in the plane cell velocity vs. cell length. (d) Examples of the three different motility modes  — sessile, transient oscillation and stable oscillation. Shown is the cell velocity over time. (e) Kymographs visualizing the pull-push-mechanism. Shown are the distributions of both myosin isoforms (upper and middle panel) and the resulting total active stress (lower panel) in a stable oscillatory state at $Pe=16000$. Parameters as in Fig. 2, except for concentration ratio of $c_{r}=1$ and strongly increased dissociation/diffusion ratio $K_{r}=10$.
The myosin isoforms are mobile within the cell and are transported with actin filaments in the retrograde actin flow. Using finite volume simulations we can trace this flow $\tilde{v}$, introduced in Eq. (10). The result can be seen in Fig. 3 as a kymograph, showing a cell that, after an initial perturbation, approaches a state of steady motility to the left. The retrograde flow is directed towards the rear end of the cell with a typical speed of $\tilde{v}_{\text{max}}\sim 10^{-2}\mu\mathrm{m}\,\mathrm{s}^{-1}$. This value is small, but not unrealistic [54].
@@ III.4 Exploring Parameter Space and Occurrence of Oscillations
We now lift the constraint of steady states, solving the full boundary value problem Eq. (10) numerically, and explore the parameter space beyond the values estimated from experiments. As in the polarization mechanism, the most interesting parameters to study are the ratio of the dissociation constants, $K_{r}$, the contractilities $P_{A}$ and $P_{B}$, and the Péclet number. These parameters again qualitatively determine the motility mode of the system. In the following, we will keep the contractility values, but consider a large $K_{r}=10$ instead of the low $K_{r}=0.2$ from before. In this parameter regime, we find additional modes of motility, in particular the possibility of oscillations, as shown in Fig. 4. For small Péclet number Pe we again find stable sessile steady states. Increasing Pe leads to the onset of transient oscillations, whose amplitudes decay towards the sessile steady state in the regime $6000\lesssim Pe\lesssim 11500$. Beyond a critical value of $Pe_{\text{crit}}\approx 11500$, these oscillations prevail and become stable. This behavior can be traced back to a Hopf bifurcation as shown in Fig. 4(a)-(c). Examples for the three different modes of motility — sessile, transient oscillation and stable oscillation — are shown in Fig. 4 (d).
@@
How can the mechanism of the oscillation be understood? One has to note that for $K_{r}>0$, one myosin isoform – namely NMII B – is more diffusive and, because of our choice of parameters here, also more contractile. If we initially assume a spatial separation of both isoforms, the stronger isoform B then induces a larger — and hence dominating — contractile stress that induces an actin flow towards the position where isoform B is the most aggregated. Myosin motors of isoform A are advected along the retrograde actin flow. As an effect of volume exclusion, isoform A subsequently displaces the more diffusive isoform B from its initial position and effectively ”pushes” it to the other side of the cell. From this point on the pull-and-push cycle repeats. A schematic sketch of the internal myosin dynamics during one oscillation cycle is depicted in Fig. 5.